OpenStudy (anonymous):

A triple scoop of ice cream is to be picked from a group of 5 flavors: vanilla, chocolate, strawberry, coffee, and peach. What is the probability that both vanilla and chocolate will be on the same cone if each flavor cannot be picked more than once on each cone? A. 1/5 B. 3/10 C. 2/5 D. 3/5 E. 2/3 I know that it is B. 3/10 but could someone tell me how please? :3 thanks :)

OpenStudy (anonymous):

is there a more mathematical way of explaining this? like for other problems if there's 3 styles and 5 colors there would be 15 combos...

OpenStudy (loser66):

@phi @sourwing please

OpenStudy (anonymous):

well, so you have vanillia and chocolate, and another flavor (which you have 3 choices to pick) so, (1/5) (1/4) 3 = 3/10

OpenStudy (anonymous):

opps, I forgot to multiply by 2

OpenStudy (anonymous):

(1/5) (1/4) (2) (3) = 3/10

OpenStudy (anonymous):

where did the numbers come from? :3

OpenStudy (anonymous):

3 came from the fact that it is a triple scoop. Once you have chosen chocolate and vanilla, and since you can't choose a same flavor, you have choices left. I.e Strawberry, coffee, and peach

OpenStudy (anonymous):

you have 3 choice left for the last scoop

OpenStudy (anonymous):

so... 2 is from vanilla and chocolate 3 is from triple scoop... what about the 1/5 and 1/4 ?

OpenStudy (anonymous):

actually, 2(1/5)(1/4) goes together. you have chocolate to be chosen for the first scoop, (1/5), but only 4 choices for the second scoop, out of which 1 is vanilla, (1/4). you multiply by 2 because you could have chosen vanilla first, then chocolate (VC or CV)

OpenStudy (anonymous):

ohhhh okay :) thanks!